Happy Times currently has an all-cash policy. It is considering making a change in the credit policy by going to terms of net 30 days. Based on the following information, what do you recommend? The required return is 0.95 percent per month.

Price per unit Cost per unit Unit sales per month Current policy $295 $230 1,105 New policy $302 $234 1,125
Answer #1

The cash flow from the old policy is: = (295 – 230)*1105 = $71825

And the cash flow from the new policy will be: = (302 – 234)*1125 = $76500

The incremental cash flow, which is perpetuity, is the difference between the old policy cash flows and the new policy cash flows, so: Incremental cash flow = 76500 - 71825 = $4675

Cost of new policy = – [PQ + Q (v ′ – v) + v ′ (Q ′ – Q)]

In this cost equation, we need to account for the increased variable cost for all units produced.

This includes the units we already sell, plus the increased variable costs for the incremental units.

So, the NPV of switching credit policies is:

NPV = –[(295)(1105) + (1105)($234 – 230) + ($234)(1125 – 1105)] + ($4675/0.0095)

NPV = $157030.3